OpenStudy (anonymous):

x2 + y2 = 4 center (, ) radius =

OpenStudy (anonymous):

center(0,0) r=2

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

general form is \((x-h)^2+(y-k)^2=r^2\) for a circle with center \((h,k)\) and radius \(r\) put your example in this form by writing \((x-0)^2+(y-0)^2=2^2\) and then you can read off the answer, center is \((0,0)\) radius is 2

OpenStudy (anonymous):

Thanks you both,i think I have it now